OpenStudy (anonymous):

help please! :( 17+4p=5p-(4+p)

OpenStudy (anonymous):

\[17 + 4p = 5p-4-p\] \[4p-5p+p = -17-4\] \[0p=-21\] There are no solutions.

OpenStudy (anonymous):

I am getting no solution too

OpenStudy (anonymous):

That's good then.

OpenStudy (anonymous):

i think you have to find what p equals???

OpenStudy (anonymous):

p does not equal because there is no solution

OpenStudy (anonymous):

Not all equations have a solution and are undefined